OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -5 - 5 cos θ Origin only x-axis only y-axis only No symmetry

OpenStudy (johnweldon1993):

Can you draw it for us? *The picture makes it obvious*

OpenStudy (johnweldon1993):

Alright well here's a hint...you have a cos function (an even function)....to test if the graph is symmetric about the x-axis...you replace θ in the equation with -θ and see if the resulting equation is the same r = -5 - 5 cos θ after replacing θ with -θ ... r = -5 - 5 cos(-θ) Now we know that cos is an even function so that cos(-θ) = cos(θ) so this means that the resulting equation would be r = -5 - 5 cos θ just like your original equation...so your graph would be symmetric about the x-axis

OpenStudy (johnweldon1993):

Just in case that's confusing....Here's a good rule of thumb to remember about polar functions and symmetry... A polar curve is symmetric about the x-axis if replacing θ by -θ in its equation produces an equivalent equation, symmetric about the y-axis if replacing θ by pi-θ in its equation produces an equivalent equation, and symmetric about the origin if replacing r by -r in its equation produces an equivalent equation.

OpenStudy (anonymous):

thank you @johnweldon1993

OpenStudy (johnweldon1993):

No problem!